# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ III.1.1 Excitations in the unperturbed 1D HM
Figure 1: Excitation from the ground state at half filling in the 1D HM for $U/t=6$. (a) $A_{\rm GS}(k,\omega)t$, (b) $S(k,\omega)t/3$, (c) $N(k,\omega)t/4$. Gaussian broadening is used with a standard deviation of $0.1t$. (d)–(f) Dispersion relations obtained using the Bethe ansatz. (d) Holon modes (brown curves), spinon modes (blue curves), and holon-spinon continua (light-yellow regions) in electronic excitation. (e) 2-spinon continuum (light-yellow region) and lower edge of the continuum (blue curve) in spin excitation. (f) 2-holon continuum (light-yellow region) in charge excitation. The solid green lines indicate $\omega=0$.
@@
In the lower band ($\omega<0$) of the electronic excitation, the fan-shaped dispersion relation is identified as the holon modes [brown curves for $\omega<0$ in Fig. 1(d)], and the low-$|\omega|$ mode for $0\leq k<\pi/2$ is identified as the spinon mode [blue curve for $\omega<0$ in Fig. 1(d)]. The dispersion relations in the upper band ($\omega>0$) are obtained using the particle–hole transformation [Appendix A.1; Eq. (169)]:
@@ III.1.2 $\eta$-SU(2) symmetry
Each eigenstate of the Hamiltonian with or without the chemical-potential term or Zeeman term in the HM, PAM, and KLM can have specific values of $S$, $S^{z}$, $\eta$, and $\eta^{z}$ Essler ; TakahashiBook ; SO4 , where $\eta$ and $\eta^{z}$ denote the total $\eta$ quantum number and eigenvalue of ${\hat{\eta}}^{z}$, respectively. The $\eta$ operators are defined as
@@ III.1.3 Excitations in the unperturbed 2D HM
Figure 2: Excitation from the ground state at half filling in the 2D HM for $U/t=9$. (a) $A_{\rm GS}({\bm{k}},\omega)t$. (b) Lowest spin-excitation energies in a $4\times 4$-site cluster (red diamonds). The blue curve indicates the dispersion relation of the spin-wave mode extracted from (c). (c) $\frac{1}{2\pi}{\rm Im}\chi({\bm{k}},\omega)t$ for $\omega>0$. Gaussian broadening is used with a standard deviation of $0.1t$ in (a) and (c). The solid green lines indicate $\omega=0$.
@@ III.2.1 Overall spectral features of unperturbed spin-gapped Mott and Kondo insulators
Figure 3: Electronic excitation from the ground state at half filling in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a), (g)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(b), (h)], 1D KLM for $J_{\rm K}/t=4$ [(c), (i)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d), (j)], 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(e), (k)], and 2D KLM for $J_{\rm K}/t=8$ [(f), (l)]. (a)–(f) $A_{\rm GS}(k_{\parallel},0,\omega)t+A_{\rm GS}({k_{\parallel}},\pi,\omega)t$ [(a)], $A_{\rm GS}({\bm{k}},\omega)t$ [(b), (c), (f)], ${\bar{A}}_{\rm GS}({\bm{k}_{\parallel}},0,\omega)t+{\bar{A}}_{\rm GS}({\bm{k}_{\parallel}},\pi,\omega)t$ [(d)], and ${\bar{A}}_{\rm GS}({\bm{k}},\omega)t$ [(e)]. Gaussian broadening is used with standard deviations of $0.16t$ [(a), (b)], $0.12t$ [(c)], $0.32t$ [(d), (e)], and $0.24t$ [(f)]. (g)–(l) Dispersion relations of electronic excitations in the effective theory; $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}$ (solid blue curves), $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}$ (solid brown curves), $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}$ (solid purple curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}$ (solid orange curves), $\omega=\varepsilon^{{\cal A}{\cal T}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions for $\omega>0$), and $\omega=-\varepsilon^{{\cal R}{\cal T}}_{-{\bm{k}}_{\parallel};-{\bm{p}}_{\parallel}}$ (light-yellow regions for $\omega<0$). The solid green lines indicate $\omega=0$.
Figure 4: Spin excitation from the ground state at half filling in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a), (g)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(b), (h)], 1D KLM for $J_{\rm K}/t=4$ [(c), (i)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d), (j)], 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(e), (k)], and 2D KLM for $J_{\rm K}/t=8$ [(f), (l)]. (a)–(c) $S(k_{\parallel},\pi,\omega)t/3$ [(a)] and $S(k,\omega)t/3$ [(b), (c)]. Gaussian broadening is used with standard deviations of $0.16t$ [(a), (b)] and $0.12t$ [(c)]. (d)–(f) Lowest spin-excitation energies for $N_{\rm u}=\sqrt{10}\times\sqrt{10}$ (orange squares) [(f)], $3\times 3$ (magenta circles) [(f)], and $\sqrt{8}\times\sqrt{8}$ (red diamonds) [(d)–(f)]. The brown curves indicate the least-squares fitting in the form of $\omega=Jd\gamma_{{\bm{k}}_{\parallel}}+\Delta E$. (g)–(l) Dispersion relations of spin excitations in the effective theory; $\omega=e^{\cal T}_{{\bm{k}}_{\parallel}}$ (solid blue curves) and $\omega=e^{{\cal A}{\cal R}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions). The solid green lines indicate $\omega=0$.
Figure 5: Charge excitation from the ground state at half filling in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a), (b)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(d), (e)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(c)], and 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(f)]. (a), (d) $N(k_{\parallel},\pi,\omega)t/4$ [(a)] and $N(k,\omega)t/4$ [(d)]. Gaussian broadening is used with a standard deviation of $0.16t$. (b), (c), (e), (f) Dispersion relations of charge excitations in the effective theory; $\omega=e^{{\cal C}^{0}}_{{\bm{k}}_{\parallel}}$ (solid orange curves) and $\omega=e^{{\cal A}{\cal R}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions). The solid green lines indicate $\omega=0$.
@@
High-energy state of $N_{\rm e}=1$ ${\bar{\cal R}}_{\sigma}$ ${\rm A}_{\sigma}$ ${\rm R}^{+}_{\sigma} -$
@@
High-energy state of $N_{\rm e}=3$ ${\bar{\cal A}}_{\sigma}$ ${\rm G}_{\sigma}$ ${\rm P}^{+}_{\sigma} -$
@@
State of $\eta=1$ and $\eta^{z}=-1$ ${\cal C}^{-1}$ $\rm V$} $\rm V$} -
State of $\eta=1$ and $\eta^{z}=0$ ${\cal C}^{0}$ $\rm D$$^{-}$ $\rm D$$^{-} -$
State of $\eta=1$ and $\eta^{z}=1$ ${\cal C}^{1}$ $\rm Z$} $\rm Z$} -
@@
with $k_{\parallel 1}=k_{x}$ and $k_{\parallel 2}=k_{y}$ in $d$ dimensions. The effective nearest-neighbor hopping parameters $t_{\rm eff}^{X}$ of the boson and fermion $X$ are $\mathcal{O}(t^{2})$ and $\mathcal{O}(t)$, respectively, and the effective 3-unit-cell hopping parameter $t_{3{\rm uc}}^{X}$ of the fermion $X$ is $\mathcal{O}(t^{2})$ [Eqs. (248)–(252); Appendix B] KohnoHubLadder ; KohnoKLM ; KohnoTinduced ; TsunetsuguRMP .
@@
$\times\sum_{i\neq j}^{N_{\rm u}}e^{i({\bm{k}}_{\parallel}-{\bm{p}}_{\parallel})\cdot{\bm{r}}_{i}}e^{i{\bm{p}}_{\parallel}\cdot{\bm{r}}_{j}}|X\rangle_{i}|Y\rangle_{j}\prod_{l(\neq i,j)}^{N_{\rm u}}|{\cal G}\rangle_{l};$ (40)
@@ IV.2.1 Hole-doping-induced states
where $e^{\rm spin}_{\bm{p}}$ denotes the excitation energy of the spin excitation with a momentum $\bm{p}$} at half filling, and $\mu_{-}$ denotes the $\omega$ value at the top of the lower band:
@@
In hole-doped Mott insulators of the HMs, because the lower band in the momentum regime inside the Fermi sea is essentially entirely filled with electrons, the electronic states emerge in the momentum regime primarily outside the Fermi sea in the electron-addition spectrum.
In the 1D and 2D HMs, the emergence of electronic modes by hole doping in accordance with the above selection rules Kohno1DHub ; Kohno2DHub ; KohnoRPP is confirmed as follows: by representing the excitation energy of the spin mode as $e^{\cal T}_{\bm{k}}$ [Figs. 1(b), 1(e), 2(b), and 2(c)] and noting that the Fermi momentum of the hole-doped system is ${\bm{k}}^{-}_{\rm F}=\frac{\bm{\pi}}{2}$ in the small-doping limit [Figs. 1(a), 1(d), and 2(a)], according to Eq. (43), an electronic mode can emerge from the top of the lower band ($\mu_{-}$), exhibiting the dispersion relation of
@@
Figure 6: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/N_{\rm s}$ in the 1D HM for $U/t=6$. (a)–(c) $A_{X}(k,\omega)t$ for $|X\rangle$ of the hole-doped ground state at $\delta=0.05$ ($\eta/N_{\rm s}=0.025$ and $\eta^{z}=-\eta$) [(a)], electron-doped ground state at $\delta=-0.05$ ($\eta/N_{\rm s}=0.025$ and $\eta^{z}=\eta$) [(b)], and lowest-energy state of $\eta/N_{\rm s}=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c)]. Gaussian broadening is used with a standard deviation of $0.1t$. (d)–(f) Dispersion relations obtained using the Bethe ansatz for $\eta/N_{\rm s}\rightarrow+0$ in the small hole-doping limit [(d)], in the small electron-doping limit [(e)], and at half filling [(f)]. The dispersion relations of the emergent modes are $\omega=e^{\cal T}_{k-\frac{\pi}{2}}+\mu_{-}$ for $k>\frac{\pi}{2}$ [red curves in (d) and (f)] and $\omega=-e^{\cal T}_{-k+\frac{\pi}{2}}+\mu_{+}$ for $k<\frac{\pi}{2}$ [magenta curves in (e) and (f)], where $e^{\cal T}_{\bm{k}}$ denotes the excitation energy of the spin mode. The dispersion relations of the other modes and continua are the same as those in Fig. 1(d) with $\mu$ adjusted such that $\mu_{-}=0$ [(d)], $\mu_{+}=0$ [(e)], and $\mu=0$ [(f)]. The solid green lines indicate $\omega=0$.
Figure 7: Electronic excitation from the lowest-energy charge-perturbed states of $\eta/N_{\rm s}=0.025$ and $\eta^{z}/N_{\rm s}=-\delta/2$ in the 2D HM for $U/t=9$. $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ [(a)], electron-doped ground state at $\delta=-0.05$ [(b)], and lowest-energy state of $\eta/N_{\rm s}=0.025$ at $\delta=0$ [(c)]. Gaussian broadening is used with a standard deviation of $0.1t$. The red arrows indicate $\omega=e^{\cal T}_{\frac{\bm{\pi}}{2}}$ [(a)] and $\omega=e^{\cal T}_{\frac{\bm{\pi}}{2}}+\mu_{\rm h}$ [(c)]; magenta arrows indicate $\omega=-e^{\cal T}_{\frac{\bm{\pi}}{2}}$ [(b)] and $\omega=-e^{\cal T}_{\frac{\bm{\pi}}{2}}+\mu_{\rm e}$ [(c)], where $e^{\cal T}_{\frac{\bm{\pi}}{2}}$ denotes the excitation energy of the spin mode at ${\bm{k}}=\frac{\bm{\pi}}{2}$, obtained using exact diagonalization in a $4\times 4$-site cluster; $\mu_{\rm h}$ and $\mu_{\rm e}$ represent the chemical potential at $\delta=0.05$ and 0.05-0.05, respectively. The solid green lines indicate $\omega=0$.

### v2

@@ III.1.1 Excitations in the unperturbed 1D HM
Figure 1: Excitation from the ground state at half filling in the 1D HM. (a) $A_{\rm GS}(k,\omega)t$, (b) $S(k,\omega)t/3$, (c) $N(k,\omega)t/4$. (d)–(f) Dispersion relations obtained using the Bethe ansatz. (d) Holon modes (brown curves), spinon modes (blue curves), and holon-spinon continua (light-yellow regions) in electronic excitation. (e) Two-spinon continuum (light-yellow region) and lower edge of the continuum (blue curve) in spin excitation. (f) Two-holon continuum (light-yellow region) in charge excitation. The solid green lines indicate $\omega=0$.
@@
In the lower band ($\omega<0$) of the electronic excitation, the fan-shaped dispersion relation is identified as the holon modes [brown curves for $\omega<0$ in Fig. 1(d)], and the low-$|\omega|$ mode for $0\leq k<\pi/2$ is identified as the spinon mode [blue curve for $\omega<0$ in Fig. 1(d)]. The dispersion relations in the upper band ($\omega>0$) are obtained using the particle–hole transformation [Appendix A.1; Eq. (171)],
@@ III.1.2 $\eta$-SU(2) symmetry
Each eigenstate of the Hamiltonian, with or without a chemical potential $\mu$ or a magnetic field $H$, can have definite values of $S$, $S^{z}$, $\eta$, and $\eta^{z}$ Essler ; TakahashiBook ; SO4 , where $\eta$ and $\eta^{z}$ denote the total $\eta$ quantum number and eigenvalue of ${\hat{\eta}}^{z}$, respectively. The $\eta$ operators are defined as
@@ III.1.3 Excitations in the unperturbed 2D HM
Figure 2: Excitation from the ground state at half filling in the 2D HM. (a) $A_{\rm GS}({\bm{k}},\omega)t$. (b) Lowest spin-excitation energies in a $4\times 4$-site cluster (red diamonds). The blue curve indicates the dispersion relation of the spin-wave mode extracted from (c). (c) $\frac{1}{2\pi}{\rm Im}\chi({\bm{k}},\omega)t$ for $\omega>0$. The solid green lines indicate $\omega=0$.
@@ III.2.1 Overall spectral features of unperturbed spin-gapped Mott and Kondo insulators
Figure 3: Electronic excitation from the ground state at half filling in the ladder HM [(a), (g)], 1D PAM [(b), (h)], 1D KLM [(c), (i)], bilayer HM [(d), (j)], 2D PAM [(e), (k)], and 2D KLM [(f), (l)]. (a)–(f) $A_{\rm GS}(k_{\parallel},0,\omega)t+A_{\rm GS}({k_{\parallel}},\pi,\omega)t$ [(a)], $A_{\rm GS}({\bm{k}},\omega)t$ [(b), (c), (f)], ${\bar{A}}_{\rm GS}({\bm{k}_{\parallel}},0,\omega)t+{\bar{A}}_{\rm GS}({\bm{k}_{\parallel}},\pi,\omega)t$ [(d)], and ${\bar{A}}_{\rm GS}({\bm{k}},\omega)t$ [(e)]. (g)–(l) Dispersion relations of electronic excitations in the effective theory; $\omega=\varepsilon^{\cal A}_{{\bm{k}}_{\parallel}}$ (solid blue curves), $\omega=-\varepsilon^{\cal R}_{-{\bm{k}}_{\parallel}}$ (solid brown curves), $\omega=\varepsilon^{\bar{\cal A}}_{{\bm{k}}_{\parallel}}$ (solid purple curves), $\omega=-\varepsilon^{\bar{\cal R}}_{-{\bm{k}}_{\parallel}}$ (solid orange curves), $\omega=\varepsilon^{{\cal A}{\cal T}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions for $\omega>0$), and $\omega=-\varepsilon^{{\cal R}{\cal T}}_{-{\bm{k}}_{\parallel};-{\bm{p}}_{\parallel}}$ (light-yellow regions for $\omega<0$). The solid green lines indicate $\omega=0$.
Figure 4: Spin excitation from the ground state at half filling in the ladder HM [(a), (g)], 1D PAM [(b), (h)], 1D KLM [(c), (i)], bilayer HM [(d), (j)], 2D PAM [(e), (k)], and 2D KLM [(f), (l)]. (a)–(c) $S(k_{\parallel},\pi,\omega)t/3$ [(a)] and $S(k,\omega)t/3$ [(b), (c)]. (d)–(f) Lowest spin-excitation energies for $N_{\rm u}=\sqrt{10}\times\sqrt{10}$ (orange squares) [(f)], $3\times 3$ (magenta circles) [(f)], and $\sqrt{8}\times\sqrt{8}$ (red diamonds) [(d)–(f)]. The brown curves indicate the least-squares fitting in the form of $\omega=Jd\gamma_{{\bm{k}}_{\parallel}}+\Delta E$. (g)–(l) Dispersion relations of spin excitations in the effective theory; $\omega=e^{\cal T}_{{\bm{k}}_{\parallel}}$ (solid blue curves) and $\omega=e^{{\cal A}{\cal R}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions). The solid green lines indicate $\omega=0$.
Figure 5: Charge excitation from the ground state at half filling in the ladder HM [(a), (b)], 1D PAM [(d), (e)], bilayer HM [(c)], and 2D PAM [(f)]. (a), (d) $N(k_{\parallel},\pi,\omega)t/4$ [(a)] and $N(k,\omega)t/4$ [(d)]. (b), (c), (e), (f) Dispersion relations of charge excitations in the effective theory; $\omega=e^{{\cal C}^{0}}_{{\bm{k}}_{\parallel}}$ (solid orange curves) and $\omega=e^{{\cal A}{\cal R}}_{{\bm{k}}_{\parallel};{\bm{p}}_{\parallel}}$ (light-yellow regions). The solid green lines indicate $\omega=0$.
@@
High-energy state of $N_{\rm e}=1$ ${\bar{\cal R}}_{\sigma}$ ${\rm A}_{\sigma}$ ${\rm R}^{+}_{\sigma}$
@@
High-energy state of $N_{\rm e}=3$ ${\bar{\cal A}}_{\sigma}$ ${\rm G}_{\sigma}$ ${\rm P}^{+}_{\sigma}$
@@
State of $\eta=1$ and $\eta^{z}=-1$ ${\cal C}^{-1}$ $\rm V$} $\rm V$}
State of $\eta=1$ and $\eta^{z}=0$ ${\cal C}^{0}$ $\rm D$$^{-}$ $\rm D$$^{-}$
State of $\eta=1$ and $\eta^{z}=1$ ${\cal C}^{1}$ $\rm Z$} $\rm Z$}
@@
with $k_{\parallel 1}=k_{x}$ and $k_{\parallel 2}=k_{y}$ in $d$ dimensions. The effective nearest-neighbor hopping parameters $t_{\rm eff}^{X}$ of the boson and fermion $X$ are $\mathcal{O}(t^{2})$ and $\mathcal{O}(t)$, respectively, and the effective three-unit-cell hopping parameter $t_{3{\rm uc}}^{X}$ of the fermion $X$ is $\mathcal{O}(t^{2})$ [Eqs. (250)–(254); Appendix B] KohnoHubLadder ; KohnoKLM ; KohnoTinduced ; TsunetsuguRMP .
@@
$\times\sum_{i\neq j}e^{i({\bm{k}}_{\parallel}-{\bm{p}}_{\parallel})\cdot{\bm{r}}_{i}}e^{i{\bm{p}}_{\parallel}\cdot{\bm{r}}_{j}}|X\rangle_{i}|Y\rangle_{j}\prod_{l(\neq i,j)}|{\cal G}\rangle_{l};$ (42)
@@ IV.2.1 Hole-doping-induced states
where $e^{\rm spin}_{\bm{p}}$ denotes the excitation energy of the spin excitation with a momentum $\bm{p}$} at half filling, and $\mu_{-}$ denotes the $\omega$ value at the top of the lower band,
@@
In hole-doped Mott insulators of the HM, because the lower band in the momentum regime inside the Fermi sea is essentially entirely filled with electrons, the electronic states emerge in the momentum regime primarily outside the Fermi sea in the electron-addition spectrum.
In the 1D and 2D HMs, the emergence of electronic modes caused by hole doping in line with the above selection rules Kohno1DHub ; Kohno2DHub ; KohnoRPP is confirmed as follows: by representing the excitation energy of the spin mode as $e^{\cal T}_{\bm{k}}$ [Figs. 1(b), 1(e), 2(b), and 2(c)] and noting that the Fermi momentum of the hole-doped system is ${\bm{k}}^{-}_{\rm F}=\frac{\bm{\pi}}{2}$ in the small-doping limit [Figs. 1(a), 1(d), and 2(a)], according to Eq. (45), an electronic mode can emerge from the top of the lower band ($\mu_{-}$), exhibiting the dispersion relation of
@@
Figure 6: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/N_{\rm s}$ in the 1D HM. (a)–(c) $A_{X}(k,\omega)t$ for $|X\rangle$ of the hole-doped ground state at $\delta=0.05$ ($\eta/N_{\rm s}=0.025$ and $\eta^{z}=-\eta$) [(a)], electron-doped ground state at $\delta=-0.05$ ($\eta/N_{\rm s}=0.025$ and $\eta^{z}=\eta$) [(b)], and lowest-energy state of $\eta/N_{\rm s}=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c)]. (d)–(f) Dispersion relations obtained using the Bethe ansatz for $\eta/N_{\rm s}\rightarrow+0$ in the small hole-doping limit [(d)], in the small electron-doping limit [(e)], and at half filling [(f)]. The dispersion relations of the emergent modes are $\omega=e^{\cal T}_{k-\frac{\pi}{2}}+\mu_{-}$ for $k>\frac{\pi}{2}$ [red curves in (d) and (f)] and $\omega=-e^{\cal T}_{-k+\frac{\pi}{2}}+\mu_{+}$ for $k<\frac{\pi}{2}$ [magenta curves in (e) and (f)], where $e^{\cal T}_{k}$ denotes the excitation energy of the spin mode. The dispersion relations of the other modes and continua are the same as those in Fig. 1(d) with $\mu$ adjusted such that $\mu_{-}=0$ [(d)], $\mu_{+}=0$ [(e)], and $\mu=0$ [(f)]. The solid green lines indicate $\omega=0$.
Figure 7: Electronic excitation from the lowest-energy charge-perturbed states of $\eta/N_{\rm s}=0.025$ and $\eta^{z}/N_{\rm s}=-\delta/2$ in the 2D HM. $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ [(a)], electron-doped ground state at $\delta=-0.05$ [(b)], and lowest-energy state of $\eta/N_{\rm s}=0.025$ at $\delta=0$ [(c)]. The red arrows indicate $\omega=e^{\cal T}_{\frac{\bm{\pi}}{2}}$ [(a)] and $\omega=e^{\cal T}_{\frac{\bm{\pi}}{2}}+\mu_{\rm h}$ [(c)]; the magenta arrows indicate $\omega=-e^{\cal T}_{\frac{\bm{\pi}}{2}}$ [(b)] and $\omega=-e^{\cal T}_{\frac{\bm{\pi}}{2}}+\mu_{\rm e}$ [(c)], where $e^{\cal T}_{\frac{\bm{\pi}}{2}}$ denotes the excitation energy of the spin mode at ${\bm{k}}=\frac{\bm{\pi}}{2}$, obtained using exact diagonalization in a $4\times 4$-site cluster; $\mu_{\rm h}$ and $\mu_{\rm e}$ represent the chemical potential at $\delta=0.05$ and 0.05-0.05, respectively. The solid green lines indicate $\omega=0$.